### v1

# A Minkowski space embedding to understand Markov models dynamics

Abstract
To describe the behavior of Markov models as parameters are varied, I show how to embed the space of Markov models within a Minkowski space. This embedding maintains the inherent distance between different instances of the model. The coordinates of this embedding emerge from the symmetrized Kullback-Leibler divergence and are expressed in terms of thermodynamic quantities, organizing the Minkowski space into equilibrium and nonequilibrium components. With this approach, we can visualize models dynamics and gain a thermodynamic interpretation of information geometric concepts, even in far-from-equilibrium scenarios. I illustrate this approach using an analytically solvable molecular motor model.
Context and objectives
To gain insight into a system’s dynamical and thermodynamical properties, it is crucial to understand their variations across the parameter space. For Markov models, however, the space of parameters has no obvious intrinsic structure. For example, multiple combinations of parameters can lead to the same affinities and currents, creating ambiguity and arbitrariness in their nonequilibrium response [1, 2]. Analysis of parameter sensitivity also shows that most models are ’sloppy’, in the sense that they have a hierarchy of eigenvalues spanning multiple scales [3, 4].
One approach to tackle these challenges is through the lens of stochastic thermodynamics. In stochastic thermodynamics, the (thermo)dynamical properties of the system are examined as parameters are varied [6, 5]. In the linear regime, the parameter space has a natural structure based on the Fisher information metric, which is linked to equilibrium fluctuations [7, 8]. Outside the linear regime, however, the structure of the parameter space and its connections to thermodynamics has remained largely unexplored. One notable exception are dynamical equivalence classes, which provide a ’hidden’ structure in the parameter space induced by their transport properties [9]. These equivalence classes uniquely characterize the nonquilibrium conditions and parameters of the system, providing a well-defined and symmetric nonequilibrium theory [10].
Another important angle to study the space of parameters is through the lens of information geometry. In information geometry, the space of parameters (e.g., for probability distributions or Markov dynamics) is analyzed by introducing a divergence as a measure of the separation between models [12, 11]. In this context, Teoh et al. demonstrated that the symmetrized Kullback-Leibler divergence plays a special role in vizualizing probabilistic models [13, 14]. The symmetrized KL divergence induces a Minkowksi structure in the space of probabilistic models which preserves the ’distance’ between models.
This paper brings these different concepts and results together. Following the approach of Teoh et al. [14], I show that the space of Markov models possesses a Minkoswki structure induced by the symmetrized Kullback-Leibler divergence. In this Minkowski space, the embedding coordinates disantangle the system’s equilibrium and nonequilibrium components. The former is expressed in terms of kinetic activities and forces while the latter coincides with the dynamical equivalence classes and are expressed in terms of thermodynamic currents and affinities [6]. These findings provide a unified and systematic way to study nonequilibrium Markov systems and their dynamical properties.
I Thermodynamic decomposition of the space of Markov chains
Let’s consider the space $\Lambda$ of primitive Markov chains on a finite state space of size $N$. A Markov chain $P\in\Lambda$ admits a unique stationary state $\pi$ and can be depicted as a graph with $N$ states, $S$ self-transitions, and $E$ edges, each edge representing the allowed transition between states.
The space $\Lambda$ has $M=2E+S-N$ dimensions while the subspace $\Sigma$ of equilibrium systems has dimension $E+S-1$. A chain $P$ then has $E-N+1$ independent currents $J_{\alpha}$ and affinities $A_{\alpha}$ [15, 1]. Each independent current is associated with a fundamental cycle in the graph and a specific edge $\alpha$ [15, 1].
Multiple representations of the spaces $\Lambda$ and $\Sigma$ exist, e.g. using exponential families [12, 11, 16] or the rotational representation of Markov chains [17, 18]. Here I will represent the space $\Lambda$ using a thermodynamic decomposition [9]. In this decomposition, a chain $P$ is determined by an equilibrium component $E$ and by it affinities $A_{\alpha}$ as
$P=s\left[E\circ Z(A)\right]\,.$ (1)
Here the mapping $s$ transforms a positive matrix into a stochastic one:
$s[K]=\frac{1}{\rho}{\rm diag}(\beta)^{-1}\,K\,{\rm diag}(\beta)\,,$ (2)
where $\rho$ is the largest eigenvalue of $K$ and $\beta$ its largest right eigenvector. The matrix
$Z_{ij}\equiv\begin{cases}\exp\left(\pm A_{\alpha}/2\right)&\text{if $i\rightarrow j=\alpha$ in the $\pm$ direction,}\\ 1&\text{otherwise}\end{cases}$
captures the nonequilibrium (asymmetric) components of $P$. The equilibrium component $E$ of a given $P$ can be recovered as $E=s[(P\circ P^{T})^{(1/2)}]$, where $P^{(1/2)}$ denotes the Hadamard or elementwise exponentiation ($E$ is often denoted by $P^{(e)}$ in information geometry). The full space of Markov models is then obtained by varying the elements $E\in\Sigma$ and the affinities $A_{\alpha}$.
The dynamics $P$ sharing the same equilibrium component $E$ in the decomposition (1) are expected to display similar properties [9]. These dynamical equivalence classes indeed play a special role in nonequilibrium transport [20]. For example, the equilibrium and nonequilibrium fluctuations are identical within an equivalence class [19]. In addition, the currents fluctuations can be obtained from the current response curve itself and all response coefficients are fully symmetric and are expressed in terms of equilibrium coefficients [10]. The equivalence classes also capture the minimal distance, in the information-theoric sense, between models under fixed nonequilibrium conditions [2].
The decomposition (1) offers a clear thermodynamic interpretation and provides important features when studying transport properties [19, 10]. However, it still lacks a natural geometric structure or notion of distance. In the next section, I show how the space $\Lambda$ can be embedded in a Minkowski space with a clear physical interpretation.
II Intensive embedding in Minkowski space
Teoh et al. showed that the symmetrized Kullback-Leibler divergence $D_{sKL}$ offers a natural distance in the space of exponential families of probability distributions [14]. Using that distance [21], an exponential family of probability distributions $f$ with $N$ parameters can be embedded isometrically in a $N+N$-dimensional Minkowski space. That is, $D_{sKL}$ can be decomposed into $N+N$ components $S^{\pm}_{i}(f)$ such that
$D_{sKL}(f,f^{\prime})=\sum_{i=1}^{N}\left[\left(\Delta S^{+}_{i}\right)^{2}-\left(\Delta S^{-}_{i}\right)^{2}\right]\,,$
with $\Delta S_{i}^{\pm}=S^{\pm}_{i}(f)-S^{\pm}_{i}(f^{\prime})$. Using these $N+N$ coordinates, the models can be visualized as control parameters are varied, preserving the distance between probability distributions. All pairwise distances between model instances are given by the symmetrized Kullback-Leibler divergence [14].
In this section I adapt this approach to the context of Markov chains and show that the resulting coordinates have a clear physical interpretation in terms of kinetic and thermodynamical quantities.
The symmetrized Kullback-Leibler divergence between two Markov chains reads
$D_{sKL}(P||G)$ $=$ $D_{KL}(P||G)+D_{KL}(G||P)$ (3)
$=$ $\sum_{ij}\left[\pi_{i}P_{ij}-\nu_{i}G_{ij}\right]\ln\frac{P_{ij}}{G_{ij}}\geq 0\,.\quad$
Here $D_{KL}(P||G)=\sum_{ij}\pi_{i}P_{ij}\ln P_{ij}/G_{ij}$ is the Kullback-Leibler divergence and $\pi$ and $\nu$ are the stationary distributions of $P$ and $G$, respectively.
To calculate and recast the divergence (3) in a more useful form, I now use the thermodynamic decomposition $P=s[E\circ Z]$ and $G=s[E^{\prime}\circ Z^{\prime}]$. As demonstated in the appendix, this leads to
$D_{sKL}(P||G)$ $=$ $\frac{1}{2}\sum_{\alpha}(A_{\alpha}-A_{\alpha}^{\prime})(J_{\alpha}-J_{\alpha}^{\prime})$ (4)
$+$ $\frac{1}{2}\sum_{e}(X_{e}-X_{e}^{\prime})(Y_{e}-Y_{e}^{\prime})\,.$
Here $J_{\alpha}$ is the average current with corresponding affinity $A_{\alpha}$ while $Y_{e}=\pi_{i}P_{ij}+\pi_{j}P_{ji}$ is the activity along the edge $e=i\longleftrightarrow j$ and $X_{e}=\ln P_{ij}P_{ji}$. By analogy with the affinities, the quantities $X_{e}$ play the role of ’kinetic forces’ driving the activity $Y_{e}$ along the edge [22].
Not all variables $Y_{e}$, however, are independent since $\sum_{e}Y_{e}=1$. One variable $r$ can thus be eliminated as $Y_{r}=1-\sum_{e\neq r}Y_{e}$. With this substitution, Eq. (4) retains the same form but with effective forces
$\bar{X}_{e}=X_{e}-X_{r}\,.$
For notational simplicity, in the rest of the paper I assume that one of the dependent edge $e$ has been eliminated and write $X_{e}=\bar{X}_{e}$.
The divergence (4) between two models $P$ and $G$ comprises two types of contributions. The first contribution captures the nonequilibrium (or antisymmetric) aspects and is expressed in terms of the thermodynamic current and affinities, while the second contribution reflects the changes in the equilibrium (or symmetric) part of the dynamics. As I will argue below, the divergence (4) measures both the time and ’spatial’ asymmetry between two models.
Let’s now express the divergence (4) in terms of space-time coordinates. As shown in the appendix, the coordinates
$S_{\alpha}^{\pm}=\frac{1}{4}(A_{\alpha}\pm J_{\alpha})\quad{\rm and}\quad S_{e}^{\pm}=\frac{1}{4}(X_{e}\pm Y_{e})$ (5)
provide an intensive embedding in a $M+M$ Minkowski space:
$D_{sKL}(P||G)=\sum_{l=1}^{M}\left[(\Delta S^{+}_{l})^{2}-(\Delta S^{-}_{l})^{2}\right]\geq 0\,,$ (6)
where the sum runs over $l=(\alpha,e)$ with $e\neq r$.
The Minkowski space is $M+M$ dimensional, where $M=2E+S-N$ corresponds to the number of independent parameters of the Markov models. The spacelike coordinates $S^{+}$ are greater than the timelike coordinates $S^{-}$ for each variable, $[S^{+}_{l}(P)-S^{+}_{l}(G)]^{2}\geq[S^{-}_{l}(P)-S^{-}_{l}(G)]^{2}$. It is also possible to center and rescale the coordinates by factors $\lambda_{l}$, e.g. to minimize the sum of the square of the coordinates or adjust the units [14].
Using coordinates (5), the set of Markov chains can thus be vizualized as a manifold as control parameters are varied, preserving the symmetric divergence between systems and with clear thermodynamic interpretation of the coordinates. This constitues our main result.
III Kinetic and thermodynamics properties of the model manifold
I now explore the kinetic and thermodynamic properties of the embedding (6) before analyzing an example of molecular motor in Section IV. Taken together, the results below show that the symmetrized KL divergence measures both the time asymmetry (through the currents and affinities) and the spatial asymmetry (through the activities and kinetic forces) of the dynamics.
III.1 Equilibrium and iso-affinity manifolds
Let’s first consider the case where the affinities between models are identical, $A_{\alpha}=A^{\prime}_{\alpha}$. In these sub-manifolds, the separations $\Delta S_{\alpha}^{\pm}$ vanish and the divergence (4) simplifies to
$D_{sKL}(P||G)$ $=$ $\sum_{e}\left[(\Delta S^{+}_{e})^{2}-(\Delta S^{-}_{e})^{2}\right]$
$=$ $\frac{1}{2}\sum_{e}(X_{e}-X_{e}^{\prime})(Y_{e}-Y_{e}^{\prime})\,.$
The embedding coordinates and the separation between two models are directly expressed in terms of the systems’ activities and kinetic forces $X_{e}$ and $X^{\prime}_{e}$, which in turn measure the inhomogeneity (spatial asymmetry) of the dynamics across the graph.
Note, however, that the coordinates $S^{\pm}_{\alpha}$ still vary along the iso-affinity manifold even if $\Delta S^{\pm}_{\alpha}=0$. Only when $A_{\alpha}=0$ are the coordinates $S^{\pm}_{\alpha}$ independent of $S^{\pm}_{e}$ (more precisely, $S^{\pm}_{\alpha}=0$ in that case). The equilibrium manifold is thus special in the sense that it is entirely described by the coordinates $S^{\pm}_{e}$.
III.2 Dynamical equivalence classes
Dynamical equivalence classes play a special role in nonequilibrium response theory [10]. Remarkably, they emerge naturally in this framework.
A dynamical equivalence class is obtained when all the model instances share the same equilibrium component $E$, i.e. when $X_{e}=X_{e}^{\prime}$. In that case the separations $\Delta S^{\pm}_{e}$ vanish and the symmetrized KL divergence (4) simplifies to
$D_{sKL}(P||G)$ $=$ $\sum_{\alpha}\left[(\Delta S^{+}_{\alpha})^{2}-(\Delta S^{-}_{\alpha})^{2}\right]$
$=$ $\frac{1}{2}\sum_{\alpha}(A_{\alpha}-A_{\alpha}^{\prime})(J_{\alpha}-J_{\alpha}^{\prime})\,.$
The embedding coordinates and the separation between two models in an equivalence class are directly expressed in terms of the system’s thermodyamic currents and affinities, which in turn measure the time asymmetry of the dynamics.
As in the case of the iso-affinity manifolds, the coordinates $S^{\pm}_{e}$ vary along the dynamical equivalence class even if $\Delta S^{\pm}_{e}=0$. In contrast to the equilibrium case however, even when $X_{e}=0$ the coordinates $S^{\pm}_{e}=\pm Y_{e}$ still depend on $S^{\pm}_{\alpha}$. Nonetheless, the class $X_{e}=0$ can act as a distinct reference dynamics: it is obtained from $E=s[T]$, where $T$ is the adjacency matrix of $P$, and thus corresponds to Markov models derived from a ’spatially homogeneous’ equilibrium component (this will be illustrated in the molecular motor example in section IV).
III.3 Entropy production
The dissipation or entropy production $\Delta_{i}S=\sum_{\alpha}J_{\alpha}A_{\alpha}\geq 0$ plays an important role in thermodynamics. In the present framework, the entropy production emerges as the symmetrized KL divergence with respect to the equilibrium component $E=P^{(e)}$ [23]:
$\Delta_{i}S=\sum_{\alpha}J_{\alpha}\times A_{\alpha}=2\,D_{sKL}\left(P||E\right)$
since $J_{\alpha}^{\prime}=A_{\alpha}^{\prime}=0$ and $X_{e}^{\prime}=X_{e}$ as $E$ belong to the equivalence class of $P$.
The entropy production can also be obtained as $D_{sKL}(P||P^{(m)})$ where $P^{(m)}=(P+P^{*})/2$ with $P^{*}$ the time-reversed dynamics [23]. This occurs because $J_{\alpha}^{\prime}=A_{\alpha}^{\prime}=0$ and $Y_{e}^{\prime}=Y_{e}$ in that case.
Finally, the entropy production can also be expressed as the divergence between a dynamics and its time-reserval [24, 23]: $\Delta_{i}S=2D_{sKL}(P||P^{*})$, as seen from Eq. (4) and using that the currents and affinities of $P^{*}$ take opposite values compared to $P$ while having the same equilibrium components [25].
IV A molecular motor example
Consider a Markov chain representing a molecular motor with $2\ell$ states corresponding to different conformations of the protein complex. These states form a cycle of periodicity $2\ell$ corresponding to a revolution by 360° for a rotary motor or a reinitialization step for a linear motor. The motor alternates between two types of states according to the transition matrix [26]
$P=\begin{pmatrix}0&p_{1}&&&&1-p_{1}\\ 1-p_{2}&&p_{2}&&&\\ &1-p_{1}&&p_{1}&&\\ &&\ddots&0&\ddots&\\ &&&1-p_{1}&0&p_{1}\\ p_{2}&&&&1-p_{2}&0\end{pmatrix}_{2\ell\times 2\ell}.$
The matrix $P$ is doubly stochastic, so that its stationary state is given by the uniform distribution ${\bf\pi}=(1,1,\cdots,1)/2\ell$ for all parameters $(p_{1},p_{2})$. The average current $J$ and affinity $A$ take the form
$J=\frac{1}{2\ell}\left(p_{1}+p_{2}-1\right)$ (7)
and
$A=\ell\,\ln\frac{p_{1}p_{2}}{(1-p_{1})(1-p_{2})}\,.$ (8)
The system is at equilibrium when $J=A=0$ (i.e., when $p_{2}=1-p_{1}$).
The activities $Y_{e}=J_{e}^{+}+J_{e}^{-}$ along the edges read $Y_{1}=(p_{1}-p_{2}+1)/2\ell$ and $Y_{2}=(p_{2}-p_{1}+1)/2\ell$ depending if the transition is odd or even, respectively. However, only one edge is independent due to the symmetry of the model and the constraint $Y_{1}+Y_{2}=1/\ell$. Taking
$Y=Y_{1}=\frac{1}{2\ell}(p_{1}-p_{2}+1)$ (9)
as the independent activity leads to an effective force
$X=X_{1}-X_{2}=\ell\,\ln\frac{p_{1}(1-p_{2})}{p_{2}(1-p_{1})}\,.$ (10)
Using expressions (8) and (10), the model parameters can be expressed as [27]
$p_{1}=\frac{{\rm e}^{(A+X)/2\ell}}{1+{\rm e}^{(A+X)/2\ell}}\,,\quad\quad p_{2}=\frac{{\rm e}^{(A-X)/2\ell}}{1+{\rm e}^{(A-X)/2\ell}}\,.$ (11)
The iso-affinity classes are thus obtained by fixing $A$ and varying $X$ while the dynamical equivalence classes are obtained by varying the affinity $A$ at fixed value of $X$ [27]. The equilibrium manifold $A=0$ is given by the line $p_{2}=1-p_{1}$ (i.e., reversible motors) and the dynamical class $X=0$ by the line $p_{1}=p_{2}$ in parameter space (i.e., spatially homogeneous motors).
Figure 1: Current and activity response curves for the molecular motor. (a) Current response curves for different dynamical equivalence classes $X$. Each response curve is anti-symmetric and invariant under the transformation $X\rightarrow-X$ (not shown). (b) Activity response curves for different equivalence classes $A$. Each response curve is anti-symmetric (after translation) and invariant under the transformation $A\rightarrow-A$ (not shown).
Within these manifolds or equivalence classes, the response curves are well-defined (Fig. 1). The current response curves are anti-symmetric and capture fluctuation at all orders [27]. The activity response curves are also antisymmetric (after translation to the origin) and invariant under the transformation $A\rightarrow-A$. I contend that the activity response curves $Y(X)$ satisfy symmetry relations similar to $J(A)$, but I will keep their analysis for a subsequent paper.
The symmatrized KL divergence between two models $P=(p_{1},p_{2})$ and $G=(g_{1},g_{2})$ reads
$2\,D_{sKL}$ $=$ $(p_{1}-g_{1})\ln\frac{p_{1}}{g_{1}}+(p_{2}-g_{2})\ln\frac{p_{2}}{g_{2}}$
$+$ $(g_{1}-p_{1})\ln\frac{1-p_{1}}{1-g_{1}}+(g_{2}-p_{2})\ln\frac{1-p_{2}}{1-g_{2}}\,.$
Inserting the parametrization (11) into this expression leads to
$D_{sKL}(P||G)$ $=$ $\frac{1}{2}(A_{P}-A_{G})(J_{P}-J_{G})$ (12)
$+$ $\frac{1}{2}(X_{P}-X_{G})(Y_{P}-Y_{G})\,.\quad\quad$
As expected from our general result, the divergence is expressed in terms of the system’s current and affinity on the one hand, and the activity and kinetic force on the other hand.
As a result, the four embedding coordinates take the form
$S_{\alpha}^{\pm}=\frac{1}{4}(A\pm J)\,,\quad\quad S_{e}^{\pm}$ $=$ $\frac{1}{4}(X\pm Y)\,.$
In these coordinates, the molecular motor models form a two-dimensional manifold (Fig. 2) satisfying the Minkowski structure (6).
The dimensions $S^{\pm}_{e}$ vary less than $S^{\pm}_{\alpha}$ overall, indicating a direction that is less sensitive to changes in parameters (also seen from the response curves in Fig. 1). Note that the number of embedding coordinates is independent of the size $\ell$ of the motor.
Figure 2: Model manifold for the molecular motor. The molecular model manifold is embedded into $2+2$ dimensions. (a) A three-dimensional projection of the manifold, which is colored based on the current $J$. (b) The different two-dimensional projections of the manifold. The spacelike coordinates are denoted as $S^{+}$ and the timelike by $S^{-}$. For visual clarity, the embedding coordinates are scaled as $S_{\alpha}^{\pm}=(A/\lambda\pm\lambda J)/4$ and $S_{e}^{\pm}=(X/\lambda\pm\lambda Y)/4$ with $\lambda=L$. All 2-d projections have the same scale.
The entropy production is obtained as the divergence from the equilibrium component $E=P(X,0)=P^{(e)}$ [23]:
$\Delta_{i}S=J\times A=2\,D_{sKL}\left(P||E\right)\,.$
Similarly, the entropy production is also measured as the divergence with respect to $P^{(m)}=(P+P^{*})/2=(1/2)P(p_{1}-p_{2}+1,p_{2}-p_{1}+1)$ [23]. In that case $Y_{P}=Y_{P^{(m)}}$ even if $X_{P}\neq X_{P^{(m)}}$, so that $\Delta_{i}S=2D_{sKL}(P||P^{(m)})$.
V Summary and outlook
This work brings together thermodynamic and information geometric concepts, showing that thermodynamic quantities emerge naturally from the Minkoswki structure induced by the symmetrized KL divergence. While other divergences could be used, the symmetrized KL divergence bypasses the curse of dimensionality and provides the lowest embedding dimension, at least for exponential families of probability distributions [14, 28].
The embedding coordinates structure the space of Markov models in terms of equilibrium and nonequilibrium (i.e. symmetric and antisymmetric) components. The equilibrium coordinates are expressed in terms of the system’s activities and kinetic forces while the nonequilibrium coordinates coincide with the dynamical equivalence classes and are expressed in terms of the system’s currents and affinities. The symmetrized Kullback-Leibler divergence and resulting coordinates measure the time and spatial asymmetry between models.
This framework provides a systematic way to analyze the space of Markov models and to gain insights into the behavior and design of mesocopic devices. For example, the Minkowski embedding allows to explore a system’s response in the directions ’orthogonal’ to the dynamical equivalence classes, where additional symmetries are expected to appear. The embedding can be used to determine the minimum parameters needed to describe model data or to identify the set of governing parameters for model prediction and coarse graining [4, 28]. The space-time structure could also help untangle other important phenomena such as dynamical or nonequilibrium phase transitions.
Disclaimer. This paper is not intended for journal publication.
Appendix: Derivation of the embedding coordinates
The first step to obtain Eq. (4) is to calculate the log ratios $\ln P_{ij}/G_{ij}$ that appear in Eq. (3). Using the thermodynamic decomposition (1) with $P=s[E\circ Z]$ and $G=s[E^{\prime}\circ Z^{\prime}]$, the log ratios take the form
$\ln\frac{P_{ij}}{G_{ij}}=\ln\frac{\rho^{\prime}}{\rho}+\ln\left(\frac{\alpha_{j}^{\prime}}{\alpha_{i}^{\prime}}\frac{\alpha_{i}}{\alpha_{j}}\right)+\ln\frac{Z_{ij}}{Z^{\prime}_{ij}}+\ln\frac{E_{ij}}{E^{\prime}_{ij}}\,.$ (13)
The first term $\ln\rho^{\prime}/\rho$ vanishes when inserted into expression (3) since $\sum\pi_{i}P_{ij}=\sum\nu_{i}G_{ij}=1$. The second term vanishes as well when averaged over a stochastic dynamics (see for example Lemma 4.3 (iii) in ref. [16]). The third term in Eq. (13) reads
$(1/2)\sum_{\alpha}(A_{\alpha}-A_{\alpha}^{\prime})(J_{\alpha}-J_{\alpha}^{\prime})$ (14)
when inserted into (3), where I used that $\ln Z_{ij}=\pm A_{\alpha}/2$ if the transition $i\rightarrow j$ corresponds to $\alpha$ in the positive (negative) direction, and 00 otherwise.
To calculate the last term $\ln E_{ij}/E^{\prime}_{ij}$, remember that $E_{ij}=(1/\lambda)(\gamma_{i}/\gamma_{j})\sqrt{P_{ij}P_{ji}}$ where $\lambda$ and $\gamma$ are the Perron eigenvalue and eigenvector of $\sqrt{P_{ij}P_{ji}}$, and similarly for $E^{\prime}$. Here also the terms $\ln\lambda^{\prime}/\lambda$ and $\ln(\gamma^{\prime}_{j}\gamma_{i}/\gamma^{\prime}_{i}\gamma_{j})$ will vanish when averaged over $\sum\pi_{i}P_{ij}-\nu_{i}G_{ij}$. The remaining factor then becomes
$(1/2)\sum_{e}(X_{e}-X^{\prime}_{e})\left(Y_{e}-Y^{\prime}_{e}\right)\,,$ (15)
where the sum runs over the $E+S$ edges and self-transitions $e=i\longleftrightarrow j$, $Y_{e}=J_{e}^{+}+J_{e}^{-}$ is the average activity (sum of forward and backward currents) along an edge, and $X_{e}=\ln P_{ij}P_{ji}$. Note that only $E+S-1$ factors $Y_{e}$ are independent since $\sum_{e}Y_{e}=1$. Summing the different terms gives the decomposition (4).
Finally, to obtain the embedding coordinates (5), each term in expression (14) can be written as [14]
$(1/2)(A_{\alpha}-A_{\alpha}^{\prime})(J_{\alpha}-J_{\alpha}^{\prime})=(\Delta S^{+}_{\alpha})^{2}-(\Delta S^{-}_{\alpha})^{2}$
with the two coordinates
$S^{\pm}_{\alpha}=(1/4)(A_{\alpha}\pm J_{\alpha})\,.$
The same factorization applies to the terms $(X_{e}-X^{\prime}_{e})(Y_{e}-Y^{\prime}_{e})$ in Eq. (15). $\square$